OpenStudy (cakt07):

Solve the word problem. The base of a triangle with height h and area A is given by the formula b = 2A/h . What is the base of a triangle with area 17 cm2 and height 5 cm? A. 3.4 cm B. 3.8 cm C. 6.8 cm D. 11 cm

OpenStudy (anonymous):

C

OpenStudy (cakt07):

Thanks!

OpenStudy (texaschic101):

just sub your stuff into the formula... b = 2A/h A = 17 h = 5 b = 2(17) / 5 b = 34/5 b = 6.8 simple as that :)